# Julian Martinez-Iriarte, and Gabriel Montes-Rojas

### v3

@@ 6 Empirical application
We apply the estimator to the study of the effect of public-sponsored training programs. As argued in LaLonde (1995), public programs of training and employment are designed to improve participant’s productive skills, which in turn would increase their earning potential and decrease dependency on social welfare benefits. We use data from the Job Training Partnership Act (JTPA) program, that has been extensively studied in the literature. For example, see Bloom, Orr, Bell, Cave, Doolittle, Lin, and Bos (1997) for a description, and Abadie, Angrist, and Imbens (2002) for applications. The JTPA was a large publicly-funded training program that began funding in October 1983 and continued until late 1990’s. We focus on the Title II subprogram, which was offered only to individuals with “barriers to employment” (long-term use of welfare, being a high-school drop-out, 15 or more recent weeks of unemployment, limited English proficiency, phsysical or mental disability, reading proficiency below 7th grade level or an arrest record). Individuals in the randomly assigned JTPA treatment group were offered training, while those in the control group were excluded for a period of 18 months. Our interest lies in measuring the effect of a training offer and actual training on of participants’ future earnings.
@@
We use the database in Abadie, Angrist, and Imbens (2002) that contains information about adult male and female JTPA participants and non-participants. Let $Z$ denote the indicator variable for those receiving a JTPA offer. Of those offered, 60% completed the training; of those in the control group completion rate was less than 2%. The dependent variable $Y$ is the logarithm of 30 month accumulated earnings (we exclude individuals without earnings), $Z$ is a dummy variable for the JTPA offer, $D$ is the endogenous dummy variable corresponding to the JTPA training, $X$ is a set of exogenous covariates contaning individual characteristics. [^7] The variables used as detailed in the database are: sex hsorged black hispanic married wkless13 afdc age2225 age2629 age3035 age3644 age4554. The parameter of interest is the effect of JTPA training on log earnings.
Next, we estimate the effect of JTPA training on log earnings. Results for different models appear in Table 5 where only the coefficient of JTPA training is reported. Besides OLS and 2SLS, we also consider several control function methods. We present estimate results for different estimates of heteroskedasticity in the first-stage (Heter. first-stage). The first column of results, where Heter. first-stage $h=1$, does not consider a skedastic function in the first-stage and uses just the residuals $\check{V}$. The second column (Heter. first-stage: $h=V^{2}$) applies a first-stage skedastic function using a regression of $\check{V}^{2}$ on all covariates. The third column (Heter. first-stage: $h=ln\ V^{2}$) repeats a similar exercise but using a regression of $log\ \check{V}^{2}$ on all covariates. Regarding the CF model, the first row uses the estimated CF from the first-stage ($\check{V}$). The second row considers the residual and its squared value $(\widehat{V},\widehat{V}^{2})$. The third row adds the interaction with JTPA training to accommodate a given skedastic function in the structural equation $(\widehat{V},\widehat{V}D)$. The fourth row adds the interaction and the square ($\widehat{V},\widehat{V}D,\widehat{V}^{2}$). Finally, the last row considers the case $(\widehat{V},\widehat{V}D,\widehat{V}^{2},\widehat{V}^{2}D)$. Standard errors are computed by bootstrap using 200 replications.
Table 5: Effect of JTPA training on log earnings
Heter. first-stage $h=1$ $h=V^{2}$ $h=ln(V^{2})$
0.266***
OLS (0.0312)
[0.198, 0.319]
0.115**
2SLS (0.0495)
[0.022, 0.208]
@@
0.115** 0.242*** 0.263***
$\hat{V}$ (0.0495) (0.0452) (0.0336)
[0.022, 0.208] [0.122, 0.306] [0.191, 0.322]
0.147** 0.186*** 0.213***
$(\hat{V},\hat{V}^{2})$ (0.0696) (0.0549) (0.0462)
[0.019, 0.276] [0.070, 0.279] [0.128, 0.308]
0.165 0.189*** 0.234***
$(\hat{V},\hat{V}D)$ (0.104) (0.0448) (0.0379)
[-0.039, 0.381] [0.115, 0.291] [0.160, 0.315]
0.0450 0.200*** 0.227***
$(\hat{V},\hat{V}D,\hat{V}^{2})$ (0.234) (0.0555) (0.0526)
[-0.400, 0.478] [0.090, 0.316] [0.119, 0.323]
-0.604 0.180*** 0.235***
$(\hat{V},\hat{V}D,\hat{V}^{2},\hat{V}^{2}D)$ (0.463) (0.0580) (0.0559)
[-1.486, 0.356] [0.093, 0.313] [0.108, 0.345]
Standard errors in parentheses using 200 bootstrap replications; 95 % percentile confidence interval in brackets.. *** p$<$0.01, ** p$<$0.05, * p$<$0.1.
The empirical results in Table 5 show that the OLS point estimate is 0.266, while the 2SLS estimate is 0.115. Next, we consider the CF estimates. Note that the result for the first row ($\widehat{V}$) and first column ($h=1$) coincides with 2SLS estimate. This is because in this case there is no heteroskedasticity in the first-stage, and the second-stage does not consider any interactions. The presence of heteroskedasticity documented before indicates that models using interactions as CF might be more appropriate. Thus, columns $h=V^{2}$ and $h=\ln(V^{2})$ are important. For the case where $(\widehat{V},\widehat{V}D)$, the point estimate ranges from 0.189 to 0.234, depending on the first-stage correction (and discarding the model with no correction for heteroskedasticity in the first-stage). For the case $(\widehat{V},\widehat{V}D,\widehat{V}^{2})$ estimates range from 0.20 to 0.227, while for $(\widehat{V},\widehat{V}D,\widehat{V}^{2},\widehat{V}^{2}D)$ it ranges from 0.180 to 0.235. These cases show that results are robust to the inclusion of polynomial and interaction terms. Overall, the CF results show empirical evidence that the job training is effective in increasing earning, but estimates are more modest than those from simple OLS regressions.
@@ A.1 Interpretation of $\alpha_{1}$
if we assume that $E[\epsilon]=0$. This is similar to the approach taken by Abrevaya and Xu (2023), where the term $(g(1,x)-g(0,x))\epsilon$ reflects the variability/deviations from ATE for different individuals/subpopulations. As noted by those authors, the 2SLS bias, i.e. the discrepancy between the ATE and the LATE, depends on “the degree to which heteroskedasticity depends upon treatment, as well as the average error disturbance for the compliers” (p.4).
Now, consider the case where $D$ is allowed to take a continuum of values. Now, a more general modeling as in Florens, Heckman, Meghir, and Vytlacil (2008) might be useful. Consider the quantity
@@ A.2 Linear projections
    The residual $\overline{Y}=Y-L_{[Y|X]}(X)$ is orthogonal to any linear function of $X$ in the following sense: for $\eta\in\mathbb{R}^{p_{x}}$ we have
@@
    $L_{[Y^{\prime}|X]}(x)=\underbrace{X^{\prime}}_{1\times p_{x}}\underbrace{E[XX^{\prime}]^{-1}}_{p_{x}\times p_{x}}\underbrace{E[XY^{\prime}]}_{p_{x}\times p_{y}}.$
@@ A.3 Proof of Lemma 1
The 2SLS is a minimum distance estimator — see Section 3.8 in Hayashi (2011) — based on the following moment condition: $E[\overline{Z}\,\overline{Y}]=E[\overline{Z}\,\overline{D}]\alpha_{1}$. The population version is the probability limit of $\widehat{\alpha}_{1,2SLS}$, and is given by
@@
Let $V_{D}:=D-L_{[D|Z,X]}(Z,X)$ be the residuals from the linear projection of $D$ on $(Z,X)$, with $L_{[D|Z,X]}(Z,X):=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}$. Thus we can write
$L_{[D|X]}(X)$ $=L_{[Z^{\prime}\pi_{1}|X]}(X)+X^{\prime}\pi_{2}+\underbrace{L_{[V_{D}|X]}(X)}_{=0.}$
but since $L_{[V_{D}|Z,X]}(Z,X)=0$, then $L_{[V_{D}|X]}(X)=0$. Therefore, the residualized version of $D$ with respect to $X$ is
$\underbrace{\overline{D}}_{D-L_{[D|X]}(X)}$ $=\overline{Z}^{\prime}\pi_{1}+V_{D}$ (A.4)
Now we plug-in (A.4) into the bias part of (A.3). First note that
$E[\overline{Z}\,\overline{D}]$ $=E[\overline{Z}\,\overline{Z}^{\prime}]\pi_{1}+E[\overline{Z}V_{D}]$
@@
because $V_{D}$ is orthogonal to any linear function of both $X$ and $Z$, such as $\overline{Z}^{\prime}:=Z^{\prime}-L_{[Z^{\prime}|X]}(X)$. So that
@@
Now, since $g:=g(D,X)$, and $D=m(X,Z,V),$ the final expression for the asymptotic bias is
bias $=\Sigma_{Z}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(m(X,Z,V),X)\varepsilon].$
@@ A.4 Proof of Lemma 2
Under Assumptions 3, 4, and 5, we have the following conditional expectation:
$E[Y|D,X,Z]$ $=D\alpha_{1}+X^{\prime}\alpha_{2}+\left(1+\sum_{s=1}^{k_{g}}(\theta_{s}^{d}D^{s}+X^{\prime s}\theta_{s}^{x})\right)\left(\sum_{j=1}^{k_{v}}\mu_{j}\tilde{V}^{j}\right).$ (A.5)
@@
where $U:=Y-E[Y|D,W]$ and satisfies, by construction, $E[U|D,W]=0$. Now we use linear projections on $W$. By the projection of $Y$ on $W$ we have that $L_{Y|W}(W)=L_{D|W}(W)\alpha_{1}+W^{\prime}\alpha_{w}$, so that $Y-L_{Y|W}(W)=(D-L_{D|W}(W))\alpha_{1}+U.$ Since by assumption, $E[(D-L_{D|W}(W))^{2}]\neq 0$, then
$\alpha_{1}$ $=\frac{E[(D-L_{D|W}(W))(Y-L_{Y|W}(W))]}{E[(D-L_{D|W}(W))^{2}]}$
$=\frac{E[(D-L_{D|W}(W))Y]}{E[(D-L_{D|W}(W))^{2}]}.$
@@ A.5 Proof of Proposition 1
This means that a (possibly nonlinear) least squares regression of $(D-Z^{\prime}\pi_{1}-X^{\prime}\pi_{2})^{2}$ on a transformation of $Z$ and $X$ (given by either (8) or (9)) consistently estimates $\gamma$ [^8] A classic reference for nonlinear least squares regression is Jennrich (1969). A modern treatment is given in Example 5.27 of van der Vaart (1998). This is, of course, unfeasible, so we use $\check{V}^{2}$. To be more precise, $\widehat{\gamma}$ is obtained as
@@
$\frac{1}{n}\sum_{i=1}^{n}\nabla_{\gamma}h(X_{i},Z_{i};\widehat{\gamma})\left(\check{V}_{i}^{2}-h(X_{i},Z_{i};\widehat{\gamma})^{2}\right)h(X_{i},Z_{i};\widehat{\gamma})=o_{p}(n^{-1/2}),$ (A.9)
where $\nabla_{\gamma}h(x,z;\widehat{\gamma})$ is the gradient of the map $\gamma\mapsto h(x,z;\gamma)$ evaluated at $\gamma=\widehat{\gamma}$. Combining the empirical moment conditions given in (A.7) and (A.9) we can write the two-step estimation problem in a single step with $(\widehat{\pi}_{1}^{\prime},\widehat{\pi}_{2}^{\prime},\widehat{\gamma}^{\prime})^{\prime}$ simultaneously satisfying
$\frac{1}{n}\sum_{i=1}^{n}\begin{bmatrix}Z_{i}\left(D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}-X_{i}^{\prime}\widehat{\pi}_{2}\right)\\ X_{i}\left(D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}-X_{i}^{\prime}\widehat{\pi}_{2}\right)\\ \nabla_{\gamma}h(X_{i},Z_{i};\widehat{\gamma})\left((D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}-X_{i}^{\prime}\widehat{\pi}_{2})^{2}-h(X_{i},Z_{i};\widehat{\gamma})^{2}\right)h(X_{i},Z_{i};\widehat{\gamma})\end{bmatrix}=o_{p}(n^{-1/2}).$ (A.10)
Next , we do a Taylor expansion of (A.10) around the true value of $\phi:=(\pi_{1}^{\prime},\pi_{2}^{\prime},\gamma^{\prime})^{\prime}$, apply of the law of large numbers, and multiply across by $\sqrt{n}$. To keep notation simple, recall that $D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}=h(X_{i},Z_{i};\gamma)V_{i}$. This yields
$o_{p}(1)$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{bmatrix}Z_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ X_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ \nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}\end{bmatrix}$
$+E\begin{bmatrix}-Z_{i}Z_{i}^{\prime}&-Z_{i}X_{i}^{\prime}&0\\ -X_{i}Z_{i}^{\prime}&-X_{i}X_{i}^{\prime}&0\\ -2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}Z_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}&-2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}X_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}&\Sigma_{\gamma}\\ \end{bmatrix}$
$\sqrt{n}\begin{bmatrix}\widehat{\pi}_{1}-\pi_{1}\\ \widehat{\pi}_{2}-\pi_{2}\\ \widehat{\gamma}-\gamma\end{bmatrix}.$
Note that $E[-2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}Z_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}]=E[-2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}X_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}]=0$ because by Assumption 1, $E[V|X,Z]=0$. Also note that $E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}]=0$ because we assumed $E[V^{2}|X,Z]=1$. The expression for $\Sigma_{\gamma}$ is

### v4

@@ 6 Empirical application
We apply the estimator to the study of the effect of public-sponsored training programs. As argued in LaLonde95, public programs of training and employment are designed to improve participant’s productive skills, which in turn would increase their earning potential and decrease dependency on social welfare benefits. We use data from the Job Training Partnership Act (JTPA) program, that has been extensively studied in the literature. For example, see Bloometal97 for a description, and AbadieAngristImbens02 for applications. The JTPA was a large publicly-funded training program that began funding in October 1983 and continued until late 1990’s. We focus on the Title II subprogram, which was offered only to individuals with “barriers to employment” (long-term use of welfare, being a high-school drop-out, 15 or more recent weeks of unemployment, limited English proficiency, physical or mental disability, reading proficiency below 7th grade level or an arrest record). Individuals in the randomly assigned JTPA treatment group were offered training, while those in the control group were excluded for a period of 18 months. Our interest lies in measuring the effect of actual training on of participants’ future earnings.
@@
We use the database in AbadieAngristImbens02 that contains information about adult male and female JTPA participants and non-participants. Let $Z$ denote the indicator variable for those receiving a JTPA offer. Of those offered, 60% completed the training; of those in the control group completion rate was less than 2%. The dependent variable $Y$ is the logarithm of 30 month accumulated earnings (we exclude individuals without earnings), $Z$ is a dummy variable for the JTPA offer, $D$ is the endogenous dummy variable corresponding to the JTPA training, $X$ is a set of exogenous covariates containing individual characteristics. The variables used as detailed in the database are: sex hsorged black hispanic married wkless13 afdc age2225 age2629 age3035 age3644 age4554. Note that there are then many covariates that will not allow us to implement a nonparametric strategy, and this serves for motivation for our parametric implementation. The parameter of interest is the effect of JTPA training on earnings.
Table 5: Effect of JTPA training on earnings
Coef. earningsearnings $earnings>0$ $log(earnings)$
OLS $\alpha_{1}$ 2665.7*** 2052.2*** 0.2656***
(304.7) (323.9) (0.0294)
2SLS $\alpha_{1}$ 1728.8*** 1715.7*** 0.1151**
(498.6) (518.9) (0.0485)
$\pi_{1}$ 0.6264*** 0.6463*** 0.6463***
(0.0059) (0.0062) (0.0062)
@@
CF0 $\alpha_{1}$ 2831.9*** 3071.4*** 0.1652
(1065.0) (1316.3) (0.1298)
$\alpha_{\check{V}}$ 1572.5** 697.2 0.2465
(643.9) (686.5) (0.0675)
$\alpha_{\check{V}D}$ -3232.9 -4215.8 -0.1557
(2818.8) (3698.7) (0.3641)
$\pi_{1}$ 0.6264*** 0.6463*** 0.6463***
( 0.0059) (0.0062) (0.0062)
CF1 $\alpha_{1}$ 2441.9*** 2106.5*** 0.1931***
(440.0) (495.8) (0.0585)
$\alpha_{\hat{V}}$ 712.9* 352.4 0.1142***
(402.1) (655.1) (0.0342)
$\alpha_{\hat{V}D}$ -999.7* -711.7 -0.1172**
(563.1) (1126.2) (0.0448)
$\pi_{1}$ 0.6264*** 0.6463*** 0.6463***
( 0.0059) (0.0062) (0.0062)
$\gamma_{1}$ 0.2118*** 0.2061*** 0.2061***
(0.0024) (0.0028) (0.0028)
Notes: Standard errors in parentheses. For OLS and 2SLS robust standard errors. CF0: control function with no correction for first-stage skedastic function. CF1: control function with first-stage skedastic function correction. *** p<<0.01, ** p<<0.05, * p<<0.1.
[…]
The empirical results in Table 5 show that the OLS point estimates are larger than the 2SLS estimates for all models in levels and logarithms. However, the CF models have $\alpha_{1}$ estimates closer to OLS but larger than the IV-2SLS estimators. CF0 estimates have very large standard errors for $\alpha_{1}$ when compared with all other models, and the estimates of $\alpha_{\check{V}D}$ are not statistically significant. Moreover, CF1 have standard errors that are approximately half the CF0 level, and of similar magnitude of 2SLS ones.
The empirical set-up is similar to that in Example 1 for the case of the IV being independent of the error components ($Z$, JTPA offer, is random). For convenience we copy the main result here:
$\frac{Cov[Z,Y]}{Cov[Z,D]}$ $=\alpha_{1}+\delta\gamma\frac{E[V\varepsilon]}{\pi_{1}}=\alpha_{1}+2SLS\ Bias$
Note that the mentioned example suggests that the 2SLS estimate is biased if there is both first-stage ($\gamma$) and second-stage endogenous heteroskedasticity ($\delta$). For the former, the CF1 point estimates of the skedastic first-stage model reveal that $\gamma_{1}>0$ (i.e. $\gamma$ in the Example) is statistically significant. For the latter, CF1 also shows that $\alpha_{\hat{V}D}<0$ (i.e. $\delta$ in the Example) is also significant, thus indicating endogenous heteroskedasticity. If we assume that $E[V\varepsilon]>0$, i.e. individuals with higher unobservable components in the earnings equation are more likely to select into training, then the 2SLS bias is negative (provided that $\delta<0$, $\gamma>0$, $\pi_{1}>0$). As such, 2SLS might be downward biased to estimate $\alpha_{1}$. The CF1 estimates are in line with this argument.
Overall, the CF results show empirical evidence that the job training is effective in increasing earning, but estimates are more modest than those from simple OLS regressions, but larger than 2SLS estimates.
@@ A.1 Interpretation of $\alpha_{1}$
if we assume that $E[\epsilon]=0$. This is similar to the approach taken by AbrevayaXu23, where the term $(g(1,x)-g(0,x))\epsilon$ reflects the variability/deviations from ATE for different individuals/subpopulations. As noted by those authors, the 2SLS bias, i.e. the discrepancy between the ATE and the LATE, depends on “the degree to which heteroskedasticity depends upon treatment, as well as the average error disturbance for the compliers” (p.4).
Now, consider the case where $D$ is allowed to take a continuum of values. Now, a more general modeling as in FlorensHeckmanMeghirVytlacil08 might be useful. Consider the quantity
@@ A.2 Linear projections
    The residual $\overline{Y}=Y-L_{[Y|X]}(X)$ is orthogonal to any linear function of $X$ in the following sense: for $\eta\in\mathbb{R}^{d_{x}}$ we have
@@
    $L_{[Y|X]}(x)=\underbrace{x^{\prime}}_{1\times d_{x}}\underbrace{E[XX^{\prime}]^{-1}}_{d_{x}\times d_{x}}\underbrace{E[XY^{\prime}]}_{d_{x}\times p_{y}}.$
@@ A.3 Proof of Lemma 1
The 2SLS is a minimum distance estimator — see Section 3.8 in hayashi2011econometrics — based on the following moment condition: $E[\overline{Z}\,\overline{Y}]=E[\overline{Z}\,\overline{D}]\alpha_{1}$. The population version is the probability limit of $\widehat{\alpha}_{1,2SLS}$, and is given by
@@
Apply a linear projection to the first-stage:
$L_{[D|X]}(X)$ $=L_{[Z^{\prime}\pi_{1}|X]}(X)+X^{\prime}\pi_{2}+L_{[h(X,Z)V|X]}(X)$
Therefore, the residualized version is
$\overline{D}$ $=\overline{Z}^{\prime}\pi_{1}+\overline{hV}.$ (A.4)
where $h:=h(X,Z)$ for notational simplicity. Now we plug-in (A.4) into the bias part of (A.3). First note that
$E[\overline{Z}\,\overline{D}]$ $=E[\overline{Z}\,\overline{Z}^{\prime}]\pi_{1}+E[\overline{Z}\,\overline{hV}]$
@@
because, by Assumption 1, $E[V|X,Z]=0$, which implies $E[\overline{Z}\,\overline{hV}]=0.$ So that
@@
Now, since $g:=g(D,X)$, and $D=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z)V,$ the final expression for the asymptotic bias is
bias $=\Sigma_{h}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z)V,X)\varepsilon].$
@@ A.4 Proof of Lemma 2
Since $h\equiv 1$, $V$ is identified from the first-stage in (2) as $V=D-Z^{\prime}\pi_{1}-X^{\prime}\pi_{2}$. Under Assumptions 3, 4, and 5, we have the following conditional expectation:
$E[Y|D,X,Z]$ $=D\alpha_{1}+X^{\prime}\alpha_{2}+\left(1+\sum_{s=1}^{k_{g}}(\theta_{s}^{d}D^{s}+X^{\prime s}\theta_{s}^{x})\right)\left(\sum_{j=1}^{k_{v}}\mu_{j}V^{j}\right).$ (A.5)
@@
where U:=YE[Y|D,W]U:=Y-E[Y|D,W] and satisfies, by construction, $E[U|D,W]=0$. Now we use linear projections on $W$. By the projection of $Y$ on $W$ we have that $L_{Y|W}(W)=L_{[D|W]}(W)\alpha_{1}+W^{\prime}\alpha_{w}$, so that $Y-L_{Y|W}(W)=(D-L_{[D|W]}(W))\alpha_{1}+U.$ Since by assumption, $E[(D-L_{[D|W]}(W))^{2}]\neq 0$, then
$\alpha_{1}$ $=\frac{E[(D-L_{[D|W]}(W))(Y-L_{Y|W}(W))]}{E[(D-L_{[D|W]}(W))^{2}]}$
$=\frac{E[(D-L_{[D|W]}(W))Y]}{E[(D-L_{[D|W]}(W))^{2}]}.$
If $h\not\equiv 1$, then the first-stage in (2) is given by $D=Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z)V$. Consider the residual $h(X,Z)V$, and apply conditional expectations as follows:
$E[h(X,Z)V|X,Z]$ $=h(X,Z)E[V|X,Z]=0,$ (A.7)
where $E[V|X,Z]=0$ follows from Assumption 1, and
$E[h(X,Z)^{2}V^{2}|X,Z]$ $=h(X,Z)^{2}E[V^{2}|X,Z]=h(X,Z)^{2},$ (A.8)
which follows from Assumption 6. Thus, it is in principle possible to identify $h(X,Z)$ in a two-stage procedure. First, identify $h(X,Z)V$ as the residuals from the regression of $D$ on $Z$ and $X$, which is possible by (A.7). Second, following (A.8), run a regression of $h(X,Z)^{2}V^{2}$ on $X$ and $Z$ to obtain $h(X,Z)^{2}$. Finally, construct $V$ as
$V=\frac{D-Z^{\prime}\pi_{1}-X^{\prime}\pi_{2}}{h(X,Z)},$ (A.9)
which is possible since $h(X,Z)\neq 0$ almost surely. Once $V$ is identified, we can proceed as in the case where $h\equiv 1$. That is, $\alpha_{1}$ is identified by (7).
@@ A.5 Proof of Proposition 1
This means that a (possibly nonlinear) least squares regression of $(D-Z^{\prime}\pi_{1}-X^{\prime}\pi_{2})^{2}$ on a transformation of $Z$ and $X$ (given by either (8) or (9)) consistently estimates $\gamma$. [^6] A classic reference for nonlinear least squares regression is jennrich1969. A modern treatment is given in Example 5.27 of vanderVaart98. This is, of course, unfeasible, so we use $\check{V}^{2}$. To be more precise, $\widehat{\gamma}$ is obtained as
@@
$\frac{1}{n}\sum_{i=1}^{n}\nabla_{\gamma}h(X_{i},Z_{i};\widehat{\gamma})\left(\check{V}_{i}^{2}-h(X_{i},Z_{i};\widehat{\gamma})^{2}\right)h(X_{i},Z_{i};\widehat{\gamma})=0,$ (A.12)
[…]
$\frac{1}{n}\sum_{i=1}^{n}\begin{bmatrix}Z_{i}\left(D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}-X_{i}^{\prime}\widehat{\pi}_{2}\right)\\ X_{i}\left(D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}-X_{i}^{\prime}\widehat{\pi}_{2}\right)\\ \nabla_{\gamma}h(X_{i},Z_{i};\widehat{\gamma})\left((D_{i}-Z_{i}^{\prime}\widehat{\pi}_{1}-X_{i}^{\prime}\widehat{\pi}_{2})^{2}-h(X_{i},Z_{i};\widehat{\gamma})^{2}\right)h(X_{i},Z_{i};\widehat{\gamma})\end{bmatrix}=0.$ (A.13)
Define
$m_{i}(\phi):=\begin{bmatrix}Z_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ X_{i}\left(D_{i}-Z_{i}^{\prime}\pi_{1}-X_{i}^{\prime}\pi_{2}\right)\\ \nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}\end{bmatrix}$,
so that A.13 can be written as
$\frac{1}{n}\sum_{i=1}^{n}m_{i}(\widehat{\phi})=0.$ (A.14)
Next, we do a mean value expansion of the sample moment around the true value $\phi$. By a first-order Taylor expansion, which is possible by Assumption 7,
[…]
$=\frac{1}{n}\sum_{i=1}^{n}m_{i}(\phi)+\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}\right)(\widehat{\phi}-\phi)+o_{p}(\|\widehat{\phi}-\phi\|)$
$=\frac{1}{n}\sum_{i=1}^{n}m_{i}(\phi)+E\left[\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}\right](\widehat{\phi}-\phi)$
$+\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}-E\left[\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}\right]\right)(\widehat{\phi}-\phi)+o_{p}(\|\widehat{\phi}-\phi\|).$
First, we are going to show that Assumptions 2 and 7 imply that
$\frac{1}{n}\sum_{i=1}^{n}\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}-E\left[\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}\right]=o_{p}(1),$
and $E\left[\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}\right]$ is invertible. To that end, we compute the derivative.
$\frac{\partial m_{i}(\phi)}{\partial\phi^{\prime}}=\begin{bmatrix}-Z_{i}Z_{i}^{\prime}&-Z_{i}X_{i}^{\prime}&0\\ -X_{i}Z_{i}^{\prime}&-X_{i}X_{i}^{\prime}&0\\ -2\nabla_{\gamma}h_{i}V_{i}Z_{i}^{\prime}h_{i}^{2}&-2\nabla_{\gamma}h_{i}V_{i}X_{i}^{\prime}h_{i}^{2}&\Sigma_{\gamma}\end{bmatrix}$,
where $h_{i}:=h(X_{i},Z_{i};\gamma)$. Note that
$E[-2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}Z_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}]=E[-2\nabla_{\gamma}h(X_{i},Z_{i};\gamma)V_{i}X_{i}^{\prime}h(X_{i},Z_{i};\gamma)^{2}]=0$
because by Assumption 1, $E[V|X,Z]=0$. Also note that $E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}]=0$ because we assumed $E[V^{2}|X,Z]=1$. The expression for $\Sigma_{\gamma}$ is